OpenStudy (anonymous):

do x^2 - 5x +6 and x^2 -4x +4 have a common factor

OpenStudy (anonymous):

well first you have to factor them both out

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2 &- 5x &+6\\ &\uparrow &\uparrow \\ &-3-2&-3\cdot -2 \end{array}\qquad \qquad \begin{array}{cccllll} x^2 &-4x &+4\\ &\uparrow &\uparrow \\ &-2-2&-2\cdot -2 \end{array}\) what do you think?